I sent this problem in and made an error on how it has to be solved The system is to be solved by addition. 5x-3y=13 4x-3y=11 Thanks
Found 2 solutions by funmath, stanbon:
Answer by funmath(2933)

E1: 5x-3y=13
E2: 4x-3y=11
:
This is the elimination method (also known as the addition method).
Multiply the E2 by -1 and add it to E1 and the y's will be eliminated so that you can solve for x.
:
-1(4x-3y)=-1(11) --> -4x+3y=-11
:
5x-3y=13
-4x+3y=-11
__________
x+0=2
x=2
:
Substitute 2 in for x in either equation and solve for y.
4(2)-3y=11
8-3y=11
8-8-3y=11-8
-3y=3
-3y/-3=3/-3
y=-1
:
Therefore, your solution is: (x,y)=(2,-1)
Happy Calculating!!!

Answer by stanbon(75887)

is to be solved by addition.
1st: 5x-3y=13
2nd: 4x-3y=11
Multiply 2nd by -1 to get:
3rd: -4x+3y=-11
Add 3rd to 1st to get:
x=2
Substitute into 2nd to solve for "y":
4(2)-3y=11
-3y=2
y=-1
Check in 1st:
5(2)-3(-1)=13
10+3=13
13=13
Cheers,
Stan H.